OpenStudy (anonymous):

MEDALS! What is the volume of a sphere with a radius of 6 inches? 112 cubic inches 288 cubic inches 290 cubic inches 300 cubic inches

OpenStudy (ffrocks3):

V=πr2h

OpenStudy (ffrocks3):

\[V=πr2h\]

OpenStudy (ffrocks3):

Dafuq openstudy

OpenStudy (anonymous):

@Ffrocks3 lmao!

OpenStudy (haleyelizabeth2017):

The volume of a sphere is \[\frac{ 4 }{ 3 }\pi r^3\]

OpenStudy (ffrocks3):

\[ V=\pi r ^{2} h\]

OpenStudy (ffrocks3):

There we go

OpenStudy (haleyelizabeth2017):

That is for a cylinder @Ffrocks3 :) And volume is always cubed :)

OpenStudy (anonymous):

so its 300?

OpenStudy (ffrocks3):

yep sorry^^

OpenStudy (mertsj):

\[V=\frac{4}{3}\pi r^3\]

OpenStudy (haleyelizabeth2017):

Then just plug in the value of the radius and solve :) follow PEMDAS so do the exponents first :)

OpenStudy (haleyelizabeth2017):

I'm not sure what I am doing wrong but I'm not getting any of the answers. @Mertsj can you check her answer?

OpenStudy (anonymous):

290?

OpenStudy (haleyelizabeth2017):

Are you sure it is supposed to be volume and not surface area?

OpenStudy (anonymous):

@haleyelizabeth2017 its volume.

OpenStudy (haleyelizabeth2017):

Okay. When I do the volume of it, I didn't get any of the answers.

OpenStudy (anonymous):

@haleyelizabeth2017 this is a tough question!

OpenStudy (haleyelizabeth2017):

Eh, once you get it down, it's not as hard :) So just keep practicing and soon, it should be easy :)

OpenStudy (anonymous):

@haleyelizabeth2017 ok!

OpenStudy (haleyelizabeth2017):

Are the answers supposed to be \(112 \pi in^3\) \(288 \pi in^3\) \(290 \pi in^3\) \(300 \pi in^3\) Because if that is right, then I can get the correct answer :)

OpenStudy (anonymous):

yes its suppose to be that!

OpenStudy (haleyelizabeth2017):

Okay, then I can get the correct answer :P heehee what do you think it is?

OpenStudy (anonymous):

i think its 288?

OpenStudy (haleyelizabeth2017):

Great! :D ding ding ding we have a winner! :)

OpenStudy (anonymous):

@haleyelizabeth2017 hehe! Haley you are awesome! Thank you very much! Merry Christmas!

OpenStudy (haleyelizabeth2017):

You too :D